OpenStudy (anonymous):

Solve for M: M/9 + 1/6= 1/2 Help please :)

OpenStudy (anonymous):

start by making a common denominator of 18: 2M/18 + 3/18 = 9/18 Then subtract 3/18 from 9/18 2M/18 = 6/18 Now you can do a "butterfly" cross multiplication OR since the deonimators are the same, you can ignore the 18 and just have 2M=6 in which case M=3 Hope this helped!

OpenStudy (anonymous):

TY :)

OpenStudy (anonymous):

No problem :)